A snail can move 8 inches in 11 minutes. At this rate, how many feet can it move in 33 hours?

Answer by tommyt3rd(5050)

8 in = 8/12 ft
11 mun = 11/60 hr
r=(8/12)/(11/60)=(8/12)(60/11)=40/11 ft/hr
d=r*t=(40/11)*33=120 ft